OpenStudy (anonymous):

Math Help Please!!!! Evaluate f(x) = -2x^2 – 4 for x = 3.

Nnesha (nnesha):

same like last one

Nnesha (nnesha):

here x =3 so what you need to do ???

Nnesha (nnesha):

looks good but don't forget that little "4"

Nnesha (nnesha):

negative 4**

OpenStudy (anonymous):

\[f(3) = -2(3)^{2} - 4\]

Nnesha (nnesha):

yes perfect!! now solve

OpenStudy (anonymous):

f(3) = -2(3)^2 -4 f(3) = -6 ^2 – 4 f (3) = -36 – 4 f (3) = -40

Nnesha (nnesha):

hmm so you don't have to solve right

OpenStudy (anonymous):

What did I do wrong? @Nnesha

Nnesha (nnesha):

ohh i i thought that are multiple choices

OpenStudy (anonymous):

You're suppose to square first then multiply, 3^3 = 9, 9*-2

OpenStudy (anonymous):

3^2*

Nnesha (nnesha):

first take square of 9 then multiply by -2

OpenStudy (anonymous):

Yes, there is multiple choices A) -10 B) -18 C) -22 D) 10

Nnesha (nnesha):

oaky so what you get after that

OpenStudy (anonymous):

Yeah so, do it again, and square it first.

OpenStudy (anonymous):

F(3) = -2(3) ^2 – 4 -2 (9) – 4 18 – 4 14

Nnesha (nnesha):

hmm 9 times -2 = ???

OpenStudy (anonymous):

18

OpenStudy (anonymous):

- x + = -

Nnesha (nnesha):

nope negative number times positive number = negative number

OpenStudy (anonymous):

Okay so -22 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you guys!

Nnesha (nnesha):

^^^^^^^^^^^^^^

Nnesha (nnesha):

our pleasure :) :) :)